please asssit
Solve the following equation:
2^x + 3=64^x/8.

Answer by stanbon(75887)

Solve the following equation:
2^x + 3=64^x/8.
---------
2^(x+3) = (2^6)^(x/8)
-----
2^(x+3) = 2^(3x/4)
-----
Equation:
x+3 = 3x/4
-----
4x+12 = 3x
x = -12
-----------------
Cheers,
Stan H.